0.5g sample of a sulphite salt was dissolved in 20ml solution and 20ml of this solution required 10ml of 0.02M acidified permanganate solution. Hence, the % mass of sulphite in the sulphite salt is?

Grade:11

1 Answers

Suraj Prasad IIT Patna
7 years ago
The answer coming is 8%. See the solution given below.

Lets say the salt is MSO3, M is any metal

The n factor of this salt will be obviously 2 .

Now its given 20 ml of this requires 10 ml of 0.02M acidified KMno4 ( whose n factor will be 5 as its acidic medium).

So we can write 10* 0.02 *5 = 0.5*1000/E
where E is equivalent weight,
E is coming 500 g, so the molar wt. of the salt is coming as 1000 g.

% SO3 will be = (80/1000) *100 = 8 %